Identify and explain the main elements used in the common stock dividend valuation model.
Answer #1

The dividend discount model is the model used for determining the value of common stock .

The formula for this valuation method is :

Po = D1/ Re - g

where,

D1= dividend expected to be paid next year

if the dividend paid last year is $2.5, and the growth rate for dividends is 5%, then the dividend paid next year will be :

2.5*(1.05) = $3.75

Re= cost of equity, the cost of equity can be calculated by several methods,

one such method is the CAPM:

Re = Rf + beta(Rm- Rf)

G =growth rate. The dividend growth rate can be estimated by multiplying the return on equity (ROE) by the retention ratio which is the opposite of the dividend payout ratio. (1 - dividend payout ratio = retention ratio).

If the value obtained from this model is more than the current value of the stock, then the stock is undervalued.

A stock is over valued, if the value obtained form this model is less than the market price of the stock.

When there is zero growth, the price of the stock can be calculated as :

Po = D1/ Re

where, the growth rate is zero.

the required return on equity can also be calculated by the GORDON GROWTH MDOEL AS:

Re = D1/Po + G